OpenStudy (anonymous):

Two rectangles have the same width. The length of one is 1 foot longer than the width. The length of the other is 2 feet longer than the width. The larger rectangle has 5 more square feet than the smaller. What is the width of the rectangles? @tcarroll010

OpenStudy (anonymous):

For area, we have: A = L x W The smaller rectangle has: (W)(W + 1) The larger rectangle has: (W)(W + 2) The smaller rectangle has 5 less square feet: (W)(W + 1) + 5 = (W)(W + 2) Can you solve for "W" now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Start by expanding: (W)(W + 1)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Hint: (a)(a + b) use distributive property a^2 + ab Now, use that as a guide for expanding: (W)(W + 1) This is a huge hint you've been given and all you have to do is substitute since you are given all the tools necessary to do this expansion.

OpenStudy (anonymous):

is it like this(2)(5+1)

OpenStudy (anonymous):

Yes, that is similar. (2)(5 + 1) = (2)(5) + (2)(1)

OpenStudy (anonymous):

i got 12 feet i dont know if i got it correct

OpenStudy (anonymous):

Well, (2)(5+1) = 12 But you already substituted W=5, so 5 is the answer, not 12.

OpenStudy (anonymous):

(W)(W + 1) + 5 = (W)(W + 2) W^2 + W + 5 = W^2 + 2W W = 5

OpenStudy (anonymous):

wht is it

OpenStudy (anonymous):

All set now?